# Multi-messenger observations in the Einstein Telescope era: binary neutron star and black hole - neutron star mergers

### v1

@@
Multi-messenger observations in the Einstein Telescope era: binary neutron star and black hole - neutron star mergers
@@
To date, no EM counterpart has been conclusively associated with BHNS mergers, even though these events have been observed through GWs (Abbott et al., 2021a; Abac et al., 2024). Notably, during O4a in April 2024, an event named GW230529 was reported (Abac et al., 2024). The mass of the BH in this event was , placing it within the lower mass gap between NSs and BHs. Although the probability of EM emission from this event was limited (around 10%), the potential existence of systems with low-mass BHs increases the likelihood of EM counterparts (Xing et al., 2024), even in scenarios with non-spinning BHs or soft equations of state (EoS), which appear to be the most favored cases based on LVK constraints (Abbott et al., 2021b).
@@ 2.1 NSNS and BHNS populations
For all variations, we assumed a cosmic merger rate density derived by convolving a power law delay time distribution (with a minimal delay of , Mapelli & Giacobbo 2018; Safarzadeh & Berger 2019; Zevin et al. 2020) with the cosmic star formation rate from Madau & Dickinson 2014. This was normalized to align with a local rate density . This local rate density has been obtained ensuring a consistent match with the observed frequency of significant NSNS mergers in LVK observing runs up to O4a. Given the absence of NSNS detections in the latter run, this is equivalent to the NSNS local rate density obtained in Colombo et al. (2022) considering the events in the third LVK GW transient catalog (GWTC-3, Abbott et al., 2021b), with a correction factor that stems from the increased time-volume surveyed after O4a222A detailed derivation of the correction factor for the local merger rate density is provided here: https://dcc.ligo.org/LIGO-P2400022/public.
@@
For BHNS we considered the population described in Colombo et al. (2024). Specifically, we relied on the BH and NS mass distributions from the standard parameter set (model A) detailed in Broekgaarden et al. (2021). We also incorporated the fiducial metallicity-specific star formation rate density from the same study, grounded in the phenomenological model of Neijssel et al. (2019). This rate density was normalized to be consistent with the merger rate density at redshift , reflecting all the BHNS events observed until the end of O4a (Abac et al., 2024).
@@
The binary stellar evolution model behind our BHNS population adopts the prescriptions from Fryer et al. (2012)’s ‘delayed’ supernova explosion mechanism. This results in a BH mass distribution that extends into the lower mass gap, particularly with BHs having masses below . Prior to GW observations, the existence of such low-mass BHs was widely debated, largely due to X-ray binary observations in the Milky Way, which suggested a sharp cutoff at around for BH masses (Özel et al., 2010; Farr et al., 2011). However, recent GW detections indicate that the lower mass gap may not be as empty as once believed, with observations of a number of systems that likely contain a component within this mass range (Abbott et al., 2020c; Zevin et al., 2020; Abac et al., 2024; Xing et al., 2024).
As in Colombo et al. (2024), we considered two different configurations for the BH spin parameter prior to the merger: a conservative approach with for all binaries, and a more optimistic scenario with a uniform distribution in the range , which aligns with the typical spin values obtained in various simulations (Fuller & Ma, 2019; Belczynski et al., 2020; Román-Garza et al., 2021; Bavera et al., 2020, 2021, 2023). In order to compute the NS compactness, we assumed the SFHo and the DD2 EoS, as we did for the NSNS population. In what follows, we adopt the most conservative setup with non-spinning BHs and the SFHo EoS as our fiducial population.
@@ 2.2 GW signal models and parameters estimation
with the index running over the considered detectors in the case of a network, for ET and for CE, being a cutoff frequency depending on the events’ parameters, denoting the Fourier-domain GW strain projected onto the detector and the noise power spectral density (PSD) of the th interferometer. Additionally, we calculated the 90% credible area for sky localization for each detected signal. We assumed two different configurations for ET: a triangular configuration with 10 km arms (ET), locating the detector in Sardinia, and a 2L-shaped interferometer design with 15 km arms misaligned by 45 deg among them (ET2L), locating one detector in Sardinia and the other one in the Meuse-Rhine region. For both the configurations, we also considered the possibility of ET operating in a global network with either one CE with 40 km arms or two CEs one with 40 km and one with 20 km arms (ET+2CE, ET2L+2CE), located in the US. For ET, we adopted the same sensitivity curves as in Branchesi et al. 2023333The ET PSDs are publicly available at https://apps.et-gw.eu/tds/?content=3&r=18213., while for CE we used the latest publicly available official PSDs444The CE PSDs we used are publicly available at https://dcc.cosmicexplorer.org/CE-T2000017/public.. Moreover, for each detector (and each interferometer in the case of a triangle) we incorporated a 85% uncorrelated duty cycle, aligning with the standard set in Branchesi et al. 2023. Given that the considered signals can stay in band as long as at 3G instruments, resulting in improved localization capabilities especially for a single detector, we include the effect of Earth’s rotation in the reconstruction as described in Iacovelli et al. (2022b, a). The main assumptions about the interferometers are summarized in Table 1.
@@ 2.3 Ejecta properties and EM emission models
In both the populations, we assumed a jet angular structure motivated by the one of GRB 170817A (Ghirlanda et al., 2019), characterized by a uniform core with a half-opening angle of (see again Colombo et al., 2022, for more details). However, BHNS and NSNS jets may differ due to the different environments in which they form. BHNS jets may experience less self-collimation because of the lower polar region density compared to NSNS systems (Bromberg et al., 2011; Duffell et al., 2015; Lazzati & Perna, 2019; Urrutia et al., 2021; Salafia et al., 2020; Hamidani & Ioka, 2021; Gottlieb & Nakar, 2022; Salafia & Ghirlanda, 2022). Additionally, NSNS mergers produce more isotropic ejecta and stronger post-merger winds (Foucart, 2020; Kawaguchi et al., 2016; Fernández & Metzger, 2013; Just et al., 2015). For this reason, in the BHNS population we also considered a broader jet opening angle of , while keeping other parameters the same, adjusting only the jet core isotropic-equivalent energy to maintain the total jet energy constant.
@@ 2.4 Multi-messenger detection criteria
Regarding the VHE afterglow band, we used the sensitivity curves of CTA North and South555The sensitivity curves are reported here: https://www.ctao.org/for-scientists/performance/ relative to photon energies of 0.1 and 1 TeV, assuming an integration time ranging from the minimum time of the light curve (0.01 days) up to 50 hours. We considered an event as detected if at least one point in its light curve was above this sensitivity threshold. We additionally assumed a 15% duty cycle to account for weather and moon constraints and a 50% reduction in sky visibility, considering that the sub-arrays are unable to observe the sky beyond a zenith angle of 60 degrees.
@@ 3.2 Detection rates
We find that ET will detect NSNS mergers per year, with the 90th percentile of the redshift distribution of detectable events at . Among these GW detectable binaries, have also a KN that exceeds, up to redshift , the brightness thresholds at its peak. The large number of detectable events and the rapid evolution of these sources poses a significant challenge to the EM follow up infrastructure. The ability of ET to localize the GW source conditions the effectiveness of the EM search for a counterpart, by determining the number of telescope pointings needed to adequately cover the GW localization region (Branchesi et al., 2023). For this reason, in Table 2 and Figure 3 we also report the detections rates for the best localized binaries, having . The rate of GW-detectable NSNS mergers with a potentially detectable KN that are also localized to within such an accuracy decreases to yr-1, which represents a more realistically manageable number.
@@ 3.3 Variations in the NSNS and BHNS populations
In the right-hand panel of Figure 4, we show the variations in the BHNS mergers population. As in the previous case, filled and empty markers represent the SFHo and DD2 EoS, respectively, while triangles and circles indicate the assumed BH spin distribution, either or a uniform distribution between 0 and 0.5. In this case, no variations in the mass distribution were considered: this is because we find that most of the mass distribution differences in the model variations explored in Broekgaarden et al. (2021) appear at high BH masses, while the mass range relevant for EM emission changes only slightly. The only exception is for models that assume the ‘rapid’ SN model from Fryer et al. (2012), but the support for that particular model, which produces a gap below 5 in the distribution of BH masses in binaries, has weakened in the last years, especially because of GW observations of objects that fall in such mass gap (Abac et al., 2024). Therefore, we do not explore BHNS mass distributions other than the fiducial one.
@@
GW ET KN+GW GRB Afterglow+GW GRB Prompt+GW J z g Radio Optical X-rays VHE Fermi NSNS Limit 12 21 24.4 26 0.01 26 CTA ET Rate () () () () () () () () () () ET2L Rate () () () () () () () () () () ET + 2CE Rate () () () () () () () () () () ET2L + 2CE Rate () () () () () () () () () () BHNS Limit 12 21 24.4 26 0.01 26 CTA ET Rate () () () () () () () () () ET2L Rate () () () () () () () () () ET + 2CE Rate () () () () () () () () () ET2L + 2CE Rate () () () () () () () () ()
@@ Acknowledgements.
- Broekgaarden et al. (2021) Broekgaarden, F. S., Berger, E., Stevenson, S., et al. 2021, arXiv e-prints, arXiv:2112.05763
@@
- Román-Garza et al. (2021) Román-Garza, J., Bavera, S. S., Fragos, T., et al. 2021, ApJ, 912, L23
@@
- Xing et al. (2024) Xing, Z., Kalogera, V., Fragos, T., et al. 2024, arXiv e-prints, arXiv:2410.20415

### v2

# Key Words.: relativistic astrophysics – multi-messenger astronomy –gravitational waves – kilonovae – gamma-ray burst1 Introduction

@@
Multi-messenger observations in the Einstein Telescope era: binary neutron star and black hole - neutron star mergers††thanks: The data produced in this work are publicly available on Zenodo through the link https://doi.org/10.5281/zenodo.15411012. The scripts and files to reproduce the main figures in the text are publicly available at https://github.com/acolombo140/ET_MM.
@@
To date, no EM counterpart has been conclusively associated with BHNS mergers, even though these events have been observed through GWs (Abbott et al., 2021a; Abac et al., 2024). Notably, during O4a in April 2024, an event named GW230529 was reported (Abac et al., 2024). The mass of the BH in this event was , placing it within the lower mass gap between NSs and BHs. Although the probability of EM emission from this event was limited (around 10%), the potential existence of systems with low-mass BHs increases the likelihood of EM counterparts (Xing et al., 2024b), even in scenarios with non-spinning BHs or soft equations of state (EoS), which appear to be the most favored cases based on LVK constraints (Abbott et al., 2021b).
@@ 2.1 NSNS and BHNS populations
For all variations, we assumed the same cosmic merger rate density as in Colombo et al. (2022), derived by convolving a power law delay time distribution (with a minimal delay of , Mapelli & Giacobbo 2018; Safarzadeh & Berger 2019; Zevin et al. 2020) with the cosmic star formation rate from Madau & Dickinson 2014. This was normalized to align with a local rate density , obtained ensuring a consistent match with the observed frequency of significant NSNS mergers in LVK observing runs up to O4a. Given the absence of NSNS detections in the latter run, this is equivalent to the NSNS local rate density obtained in Colombo et al. (2022) (see Appendix A.1) with a correction factor that stems from the increased time-volume surveyed after O4a222A detailed derivation of the correction factor for the local merger rate density is provided here: https://dcc.ligo.org/LIGO-P2400022/public.
@@
For BHNS we considered the population described in Colombo et al. (2024). Specifically, we relied on the BH and NS mass distributions from the standard parameter set (model A) detailed in Broekgaarden et al. (2021). We also incorporated the fiducial metallicity-specific star formation rate density from the same study, grounded in the phenomenological model of Neijssel et al. (2019). As we did for the NSNS population, the rate density was normalized to be consistent with the merger rate density at redshift , correcting the value presented in Colombo et al. (2024) to reflect all the BHNS events observed until the end of O4a (Abac et al., 2024)333We assume four BHNS events with false alarm rate less than 1 in 4 years, as reported in https://emfollow.docs.ligo.org/userguide/capabilities.html.
@@
The binary stellar evolution model behind our BHNS population adopts the prescriptions from Fryer et al. (2012)’s ‘delayed’ supernova explosion mechanism. This results in a BH mass distribution that extends into the lower mass gap, particularly with BHs having masses below . Prior to GW observations, the existence of such low-mass BHs was widely debated, largely due to X-ray binary observations in the Milky Way, which suggested a sharp cutoff at around for BH masses (Özel et al., 2010; Farr et al., 2011). However, recent GW detections indicate that the lower mass gap may not be as empty as once believed, with observations of a number of systems that likely contain a component within this mass range (Abbott et al., 2020c; Zevin et al., 2020; Abac et al., 2024; Xing et al., 2024b).
As in Colombo et al. (2024), we considered two different configurations for the BH spin parameter prior to the merger: a conservative approach with for all binaries, and a more optimistic scenario with a uniform distribution in the range . These values align with the typical spin values obtained in various simulations, assuming that the helium star progenitors of BHs that form second in the binary can spin-up through tidal interactions (Fuller & Ma, 2019; Belczynski et al., 2020; Román-Garza et al., 2021a; Bavera et al., 2020, 2021, 2023). If the BH forms first, it is believed to have zero spin as a consequence of efficient angular momentum transport (Fragos & McClintock, 2015; Qin et al., 2018; Fuller & Ma, 2019; Belczynski et al., 2020).
In order to compute the NS compactness, we assumed the SFHo and the DD2 EoS, as we did for the NSNS population. In what follows, we adopt the most conservative setup with non-spinning BHs and the SFHo EoS as our fiducial population.
@@ 2.2 GW signal models and parameters estimation
with the index running over the considered detectors in the case of a network, for ET and for CE, being a cutoff frequency depending on the events’ parameters, denoting the Fourier-domain GW strain projected onto the detector and the noise power spectral density (PSD) of the th interferometer. Additionally, we calculated the 90% credible area for sky localization for each detected signal. We assumed two different configurations for ET: a triangular configuration with 10 km arms (ET), locating the detector in Sardinia, and a 2L-shaped interferometer design with 15 km arms misaligned by 45 deg among them (ET2L), locating one detector in Sardinia and the other one in the Meuse-Rhine region. For both the configurations, we also considered the possibility of ET operating in a global network with either one CE with 40 km arms or two CEs one with 40 km and one with 20 km arms (ET+2CE, ET2L+2CE), located in the US. For ET, we adopted the same sensitivity curves as in Branchesi et al. 2023444The ET PSDs are publicly available at https://apps.et-gw.eu/tds/?content=3&r=18213., while for CE we used the latest publicly available official PSDs555The CE PSDs we used are publicly available at https://dcc.cosmicexplorer.org/CE-T2000017/public.. Moreover, for each detector (and each interferometer in the case of a triangle) we incorporated a 85% uncorrelated duty cycle, aligning with the standard set in Branchesi et al. 2023. Given that the considered signals can stay in band as long as at 3G instruments, resulting in improved localization capabilities especially for a single detector, we include the effect of Earth’s rotation in the reconstruction as described in Iacovelli et al. (2022b, a). The main assumptions about the interferometers are summarized in Table 1.
@@ 2.3 Ejecta properties and EM emission models
In both the populations, we assumed a jet angular structure motivated by the one of GRB 170817A (Ghirlanda et al., 2019), characterized by a uniform core with a half-opening angle of (see again Colombo et al., 2022, for more details). However, BHNS and NSNS jets may differ due to the different environments in which they form. BHNS jets may experience less self-collimation because of the lower polar region density compared to NSNS systems (Bromberg et al., 2011; Duffell et al., 2015; Lazzati & Perna, 2019; Urrutia et al., 2021; Salafia et al., 2020; Hamidani & Ioka, 2021; Gottlieb & Nakar, 2022; Salafia & Ghirlanda, 2022). Additionally, NSNS mergers produce more isotropic ejecta and stronger post-merger winds (Foucart, 2020; Kawaguchi et al., 2016; Fernández & Metzger, 2013; Just et al., 2015). For this reason, in the BHNS population we also considered a broader jet opening angle of ,666The effect of this variation on the rates is included in the error bars in the right panel of Figure 4. The effect on the GRB properties is shown as an orange curve in the right panels of Figures 8 and 9. while keeping other parameters the same, adjusting only the jet core isotropic-equivalent energy to maintain the total jet energy constant.
@@ 2.4 Multi-messenger detection criteria
Regarding the VHE afterglow band, we used the sensitivity curves of CTA North and South777The sensitivity curves are reported here: https://www.ctao.org/for-scientists/performance/ relative to photon energies of 0.1 and 1 TeV, assuming an integration time ranging from the minimum time of the light curve (0.01 days) up to 50 hours. We considered an event as detected if at least one point in its light curve was above this sensitivity threshold. We additionally assumed a 15% duty cycle to account for weather and moon constraints and a 50% reduction in sky visibility, considering that the sub-arrays are unable to observe the sky beyond a zenith angle of 60 degrees.
@@ 3.2 Detection rates
We find that ET will detect NSNS mergers per year, with the 90th percentile of the redshift distribution of detectable events at . Among these GW detectable binaries, have also a KN that exceeds, up to redshift , the brightness thresholds at its peak. The large number of detectable events and the rapid evolution of these sources poses a significant challenge to the EM follow up infrastructure. The ability of ET to localize the GW source conditions the effectiveness of the EM search for a counterpart, by determining the number of telescope pointings needed to adequately cover the GW localization region (Branchesi et al., 2023). For this reason, in Table 2 and Figure 3 we also report the detections rates for the best localized binaries, having or . The rate of GW-detectable NSNS mergers with a potentially detectable KN that are also localized to within such an accuracy decreases to yr-1, which represents a more realistically manageable number.
@@ 3.3 Variations in the NSNS and BHNS populations
In the right-hand panel of Figure 4, we show the variations in the BHNS mergers population. As in the previous case, filled and empty markers represent the SFHo and DD2 EoS, respectively, while triangles and circles indicate the assumed BH spin distribution, either or a uniform distribution between 0 and 0.5. In this case, we did not consider variations in the BH mass distribution. While it is clear that different assumptions on the mass distribution can affect the multi-messenger detection rates, we focused on the uncertainties associated with the BH spin, the NS EoS, and the overall merger rate. These factors, as well the systematics in the modeling, tend to dominate the impact on the fraction of events producing EM emission, as shown in Broekgaarden et al. (2021); Román-Garza et al. (2021b); Biscoveanu et al. (2023); Chen et al. (2024); Xing et al. (2024a). We note, however, that the role of BH spin must be interpreted carefully. While our models include a wide range of spin assumptions to bracket potential outcomes, realistic spin distributions from isolated binary evolution are expected to favor low BH spins, with only a small fraction of systems potentially reaching high spins due to formation order reversal (e.g. Fuller & Ma, 2019; Xing et al., 2024a). Therefore, the impact of BH spin on EM detectability may be somewhat overestimated under more optimistic spin assumptions. Additionally, we acknowledge that recent observational results challenge some long-standing assumptions about the BH mass distribution. In particular, the detection of GW230529 (Abac et al., 2024) not only questions the existence of a lower-mass gap—reproduced by the ‘rapid’ SN model of Fryer et al. (2012)—but also suggests that the typical BH mass in BHNS systems may be lower than previously assumed. While this potential shift in the BH mass distribution is not explicitly modeled in our current work, it is an important direction for future studies, since it could further enhance EM detectability in the BHNS population.
@@
GW ET KN+GW GRB Afterglow+GW GRB Prompt+GW J z g Radio Optical X-rays VHE Fermi NSNS Limit 12 21 24.4 26 0.01 26 CTA ET Rate () () () () () () () () () () () () () () () () () () () () ET2L Rate () () () () () () () () () () () () () () () () () () () () ET + 2CE Rate () () () () () () () () () () () () () () () () () () () () ET2L + 2CE Rate () () () () () () () () () () () () () () () () () () () () BHNS Limit 12 21 24.4 26 0.01 26 CTA ET Rate () () () () () () () () () () () () () () () () () () ET2L Rate () () () () () () () () () () () () () () () () () () ET + 2CE Rate () () () () () () () () () () () () () () () () () () ET2L + 2CE Rate () () () () () () () () () () () () () () () () () ()
@@ 4.3 GRB Prompt
For the BHNS population, we also report a dashed orange line corresponding to 90% of binary systems that are detectable through GRB prompt emission, assuming a jet half-opening angle deg. Varying the jet core half-opening angle while keeping the total jet energy fixed primarily results in a shift of the distribution: wider jets correspond to lower on-axis values, as this quantity scales approximately as . As a result, populations with systematically larger opening angles would trace a distinct locus in the – plane compared to BNS-associated GRBs, assuming similar intrinsic emission properties.
@@ Acknowledgements.
- Broekgaarden et al. (2021) Broekgaarden, F. S., Berger, E., Neijssel, C. J., et al. 2021, MNRAS, 508, 5028
@@
- Román-Garza et al. (2021a) Román-Garza, J., Bavera, S. S., Fragos, T., et al. 2021a, ApJ, 912, L23
- Román-Garza et al. (2021b) Román-Garza, J., Bavera, S. S., Fragos, T., et al. 2021b, ApJ, 912, L23
@@
- Xing et al. (2024a) Xing, Z., Bavera, S. S., Fragos, T., et al. 2024a, A&A, 683, A144
- Xing et al. (2024b) Xing, Z., Kalogera, V., Fragos, T., et al. 2024b, arXiv e-prints, arXiv:2410.20415